OpenStudy (anonymous):

17u-24=5/9 ------- 18

sam (.sam.):

\[\begin{array}{l} \frac{1}{18} (17 u-24)=\frac{5}{9} \\ \text{Divide both sides by }\frac{1}{18}: \\ 17 u-24=10 \\ \text{Add }24\text{ both sides} \\ 17 u=34 \\ \text{Divide both sides by }17 \\ u=2 \\\end{array}\]

OpenStudy (anonymous):

\[17u-24=\frac{5}{9}?\]